# Mixing Time Bounds for the Gibbs Sampler under Isoperimetry

### v1

@@ 5.1 Implications for Lipschitz and smooth potentials
where the function $U:\mathbb{R}^{d}\to\mathbb{R}$ is known as the potential. Many previous works assume that the potential function is either Lipschitz continuous (Lehec, 2023) or smooth (see, e.g., Ascolani et al., 2024a). We show that TV continuity of the conditional distributions of $\pi$ (Definition 2.5) is a strictly weaker condition than these regularity assumptions on $U$.
@@
To establish our result, we first demonstrate that log-Lipschitz continuity and log-smoothness extend to the marginal distributions of $\pi$. The proof of the following lemma is located in Appendix A.
@@
  • If the distribution $\pi$ is $L$-log-Lipschitz continuous, then all marginal distributions of $\pi$ are also $L$-log-Lipschitz continuous.
  • If the distribution $\pi$ is $L$-log-smooth, then all marginal distributions of $\pi$ are also $L$-log-smooth.
Lemma 5.3 complements Saumard and Wellner (2014, Theorems 3.3 and 3.8), who proved that (strong) log-concavity is also inherited by the marginal distributions of $\pi$.
@@
For simplicity, we redefine the variables as $x_{-k}=s$, $y_{-k}=t$. The joint distribution can be rewritten as $\pi(\theta,s)\propto e^{-U(\theta,s)}.$ From this, the conditional distribution is expressed as
$\pi(\theta|s)=\frac{\pi(\theta,s)}{\pi(s)}=\frac{e^{-U(\theta,s)}}{\int e^{-U(\psi,s)}\mathrm{d}\psi}=\frac{e^{-U(\theta,s)}}{e^{-V(s)}},$
@@
$\leq\sqrt{L\left\lvert s-t\right\rvert}.$
@@
Suppose that the distribution $\pi$ is $L$-log-smooth. Then, the conditional distributions of $\pi$ are $(\sqrt{L},1)$-TV continuous.
@@
$\mathrm{KL}(\pi(\cdot\mid s),\pi(\cdot\mid t))=\int(U(\theta,t)-U(\theta,s))\pi(\mathrm{d}\theta\mid s)+V(s)-V(t).$
Observe that $\nabla V(s)=\int\nabla_{s}U(\theta,s)\pi(\mathrm{d}\theta\mid s).$ It follows that
$\mathrm{KL}(\pi(\cdot\mid s),\pi(\cdot\mid t))$ $=\int(U(\theta,t)-U(\theta,s))\pi(\mathrm{d}\theta\mid s)+V(s)-V(t)$
$\leq\int\left(\nabla_{s}U(\theta,s)^{T}(t-s)+\frac{L}{2}\left\lvert s-t\right\rvert^{2}\right)\pi(\mathrm{d}\theta\mid s)$
$\quad+\nabla V(t)^{T}(s-t)+\frac{L}{2}\left\lvert s-t\right\rvert^{2}$
$\leq\int(\nabla_{s}U(\theta,s)-\nabla_{t}U(\theta,t))^{T}(t-s)\pi(\mathrm{d}\theta\mid s)+L\left\lvert s-t\right\rvert^{2}$
$\leq 2L\left\lvert s-t\right\rvert^{2}.$
@@
$\leq\sqrt{L}\left\lvert s-t\right\rvert.$
@@
Consider a $L$-log-smooth target distribution $\pi$ that satisfies $\mathrm{PI}_{1}$. Then, the conductance of the systematic-scan Gibbs kernel is bounded by
$\Phi(P_{\mathrm{SS}})\gtrsim\min\left\{1,\frac{\mathtt{p}_{1}}{\sqrt{L}d}\right\}.$
@@
$\tau(\zeta,P_{\mathrm{SS}})\lesssim\max\left\{1,\frac{Ld^{2}}{\mathtt{p}_{1}^{2}}\right\}\cdot\log\left(\frac{\sqrt{\Omega}}{\zeta}\right).$
However, the mixing time bound in Corollary 4.7 appears to be loose when $\pi$ is a Gaussian distribution, as suggested by comparisons with Roberts and Sahu (1997).
@@
Consider a $L$-log-smooth target distribution $\pi$ that satisfies $\mathrm{PI}_{1}$. Then, the conductance of the random-scan Gibbs kernel is bounded by
$\Phi(P_{\mathrm{RS}})\gtrsim\frac{1}{d\cdot\log d}\cdot\min\left\{1,\frac{\mathtt{p}_{1}}{\sqrt{L}d\cdot\log d}\right\}.$
@@
$\lambda_{2}(P_{\mathrm{RS}})\gtrsim\frac{1}{d\cdot\log d}\cdot\min\left\{1,\frac{\mathtt{p}_{1}^{2}}{L(d\cdot\log d)^{2}}\right\}.$
@@
$\tau(\zeta,P_{\mathrm{RS}})\lesssim d\cdot\log d\cdot\max\left\{1,\frac{L(d\cdot\log d)^{2}}{\mathtt{p}_{1}^{2}}\right\}\cdot\log\left(\frac{\sqrt{\Omega}}{\zeta}\right).$
Unfortunately, the spectral gap bound in Corollary 4.9 is suboptimal when $\pi$ is strongly log-concave and $L$-log-smooth. In such cases, Ascolani et al. (2024a, Corollary 3.6) established that
@@
Furthermore, we believe that the lower bound on the spectral gap in Theorem 3.7 is not tight for the Gibbs sampler. This belief is supported by the case of the random walk on the $d$-dimensional hypercube, where Trevisan (2013, Chapter 5) showed that the conductance and spectral gap are of the same order, making the first bound from Theorem 3.7 loose in this setting. The Gibbs sampler shares a similar local structure to the $d$-dimensional hypercube: each update resamples a single coordinate from its exact conditional, in analogy to the single-bit flips in the hypercube walk.
@@ 5.2 Non-log-concave examples
Continuous Gaussian mixture.
For all $t\in\mathbb{R}^{d}$, let $\varphi_{t}(\mathrm{d}x)\propto\exp(-|x-t|^{2}/2)\mathrm{d}x$ denote a Gaussian distribution with mean $t$ and covariance matrix $I_{d}$. Define the mixing distribution
$\nu(\mathrm{d}t)\propto e^{-\left\lvert\left\lvert t\right\rvert^{2}-4\right\rvert}\mathrm{d}t.$
The resulting Gaussian mixture
$\pi_{1}(\mathrm{d}x)=\int\varphi_{t}(\mathrm{d}x)\nu(\mathrm{d}t)$
is log-smooth and satisfies $\mathrm{LSI}_{2}$ (Chen et al., 2021, Subsection 4.2). However, since $\nu$ is multimodal, $\pi_{1}$ is not log-concave.
Discrete Laplace mixture.
Let $m_{1},m_{2}\in\mathbb{R}^{d}$ with $m_{1}=-m_{2}=(1,1,\dots,1)$. Consider the two-component mixture of Laplace distributions
$\pi_{2}(\mathrm{d}x)\propto\left(\frac{1}{2}\cdot e^{-\left\lvert x-m_{1}\right\rvert}+\frac{1}{2}\cdot e^{-\left\lvert x-m_{2}\right\rvert}\right)\mathrm{d}x.$
This distribution is log-Lipschitz continuous, and satisfies $\mathrm{PI}_{2}$ as shown by Chafaï and Malrieu (2010); Schlichting (2019). However, $\pi_{2}$ is not log-concave.
@@
$\pi_{3}(\mathrm{d}x)\propto e^{-\left\lvert x\right\rvert+2\cos\left\lvert x\right\rvert}\mathrm{d}x.$
This distribution is log-Lipschitz continuous and satisfies $\mathrm{PI}_{1}$ (Bobkov and Houdré, 1997). However, the added perturbation prevents $\pi_{3}$ from being log-concave.
@@ 6 Conclusion
We established verifiable conductance bounds for systematic-scan and random-scan Gibbs samplers over a broad class of target distributions, particularly those satisfying a Poincaré or log-Sobolev inequality and possessing sufficiently regular conditional distributions (Theorem 4.7 and Corollary 4.9). These bounds directly provide estimates on the mixing time in total variation distance and remain valid for either log-Lipschitz or log-smooth target distributions. In doing so, we extend prior results beyond the log-concave setting (Ascolani et al., 2024a; Wadia, 2024) and present several examples.
Our work also complements recent results showing that the spectral gaps and mixing times of systematic-scan and random-scan Gibbs samplers are closely related (Gaitonde and Mossel, 2024; Chlebicka et al., 2025). While our analysis focuses on coordinate-wise updates, our results naturally extend to blocked Gibbs samplers under corresponding block-smoothness conditions, allowing for joint updates of coordinate subsets. Moreover, our work also has implications for Metropolis-within-Gibbs sampling schemes; Ascolani et al. (2024b) examine how the conductance of generic coordinate-wise schemes compares to that of exact Gibbs sampling, while Qin et al. (2025) perform a similar analysis for the spectral gap. Their results, combined with our bounds, provide a framework for studying the convergence of broader coordinate-wise sampling methods.
Regarding dimension dependence, it is clear that our bounds on the spectral gap and mixing time are not tight, particularly for Gaussian targets (Amit, 1996; Roberts and Sahu, 1997) and, more generally, for log-smooth and strongly log-concave distributions (Ascolani et al., 2024a). We attribute this to the generality of our assumptions, which might not fully exploit the underlying structure of these distributions, and to the use of total variation distance in defining the mixing times. A natural direction for future research is therefore to refine these conductance bounds for both variants of the Gibbs sampler, particularly in terms of their dependence on the dimension and warmness parameter, as studied in Chen et al. (2020). One possibility is to study the mixing time of the Gibbs sampler under alternative probability metrics, such as appropriate Wasserstein distances, as in Wang and Wu (2014). Beyond the conductance approach, one could also explore recent techniques that may yield sharper spectral gap estimates for the Gibbs sampler, such as quasi-telescoping inequalities (also known as the spectral telescope), as developed in Qin and Wang (2024).

### v2

@@ 5.1 Implications for Lipschitz and smooth potentials
where the function $U:\mathbb{R}^{d}\to\mathbb{R}$ is known as the potential. Many previous works assume that the potential function is either Lipschitz continuous (Lehec, 2023) or smooth (see, for example, Ascolani et al., 2026a).
@@
We show that TV continuity of the conditional distributions of $\pi$ (Definition 2.5) can be obtained under the above regularity assumptions on $U$. To establish this, we first demonstrate that log-Lipschitz continuity and log-smoothness are preserved under marginalization, provided that the conditional distributions of $\pi$ satisfy a uniform $L^{2}$-Poincaré inequality. The proof of the following lemma is located in Appendix A.
@@
  • Suppose that the distribution $\pi$ is $L$-log-Lipschitz continuous. Then, all marginal distributions of $\pi$ are also $L$-log-Lipschitz continuous.
  • Suppose that the distribution $\pi$ is $L$-log-smooth and for all $k\in[d]$ and $x\in\mathbb{R}^{d}$, the conditional distribution $\pi(\cdot\mid x_{-k})$ satisfies $\mathrm{PI}_{2}$ with uniform constant $\mathtt{c}_{2}$. Then, all (d1)(d-1)-dimensional marginal distributions are $L^{\prime}$-log-smooth, where
    $L^{\prime}=L+\frac{L^{2}}{\mathtt{c}_{2}}.$
Lemma 5.3 complements the classical result that (strong) log-concavity is preserved under marginalization, as shown in Prékopa (1973); Brascamp and Lieb (1976); see also Saumard and Wellner (2014) for a survey. The assumption that the conditional distributions of $\pi$ satisfy a uniform $L^{2}$-Poincaré inequality has been widely used in the study of two-scale criteria for functional inequalities (Otto and Reznikoff, 2007; Lelièvre, 2009; Grunewald et al., 2009), in covariance estimates (Menz, 2014b), and in the analysis of spin systems (see, for example, Menz, 2014a).
As a simple sanity check, the above assumption holds when the target distribution is strongly log-concave. More precisely, if $\pi$ is strongly log-concave with constant $m>0$, then each one-dimensional conditional distribution is strongly log-concave with the same constant, and therefore satisfies a uniform $L^{2}$-Poincaré inequality (see the discussion in Subsection 2.2). Indeed, the potential function of the conditional distribution $\pi(\cdot\mid x_{-k})$ is given by
$x_{k}\mapsto U(x_{k},x_{-k}),$
where $x_{-k}$ is held fixed. Strong log-concavity of $\pi$ with constant $m>0$ is equivalent to the Hessian bound $\nabla^{2}U(x)\succeq mI_{d}$, where $I_{d}$ denotes the $d$-dimensional identity matrix and $\succeq$ the Loewner order. In particular, this implies that
$\frac{\partial^{2}U}{\partial x^{2}_{k}}(x_{k},x_{-k})\geq m,$
which is exactly the condition that $\pi(\cdot\mid x_{-k})$ is strongly log-concave with constant $m$.
@@
For simplicity, we redefine the variables as $x_{-k}=s$, $y_{-k}=t$. The joint distribution can be rewritten as $\pi(\theta,s)\propto\exp({-U(\theta,s)})$. From this, the conditional distribution is expressed as
$\pi(\theta\mid s)=\frac{\pi(\theta,s)}{\pi(s)}=\frac{e^{-U(\theta,s)}}{\int e^{-U(\psi,s)}\mathrm{d}\psi}=\frac{e^{-U(\theta,s)}}{e^{-V(s)}},$
@@
$\leq\sqrt{L\left\lvert s-t\right\rvert}$$\qed$
@@
Suppose that the distribution $\pi$ is $L$-log-smooth and for all $k\in[d]$ and $x\in\mathbb{R}^{d}$, the conditional distribution $\pi(\cdot\mid x_{-k})$ satisfies $\mathrm{PI}_{2}$ with uniform constant $\mathtt{c}_{2}$. Then, the conditional distributions of $\pi$ are $(M^{\prime},1)$-TV continuous, where
$M^{\prime}=\sqrt{\frac{L}{2}+\frac{L^{2}}{4\mathtt{c}_{2}}}.$
@@
$\mathrm{KL}(\pi(\cdot\mid s),\pi(\cdot\mid t))=\int(U(\theta,t)-U(\theta,s))\pi(\mathrm{d}\theta\mid s)+V(s)-V(t).$ (17)
We now bound both differences on the right-hand side of (17). By $L$-smoothness of the function $U$,
$U(\theta,t)\leq U(\theta,s)+\nabla_{s}U(\theta,s)^{T}(t-s)+\frac{L}{2}\left\lvert s-t\right\rvert^{2},$ (18)
where $\nabla_{s}U(\theta,s)$ represents the gradient of $U(\theta,s)$ with respect to the coordinates of $s$. Moreover, by Lemma 5.3, the function $V$ is $L^{\prime}$-smooth with $L^{\prime}=L+L^{2}/\mathtt{c}_{2}$. Therefore,
$V(t)\geq V(s)+\nabla V(s)^{T}(t-s)-\frac{L^{\prime}}{2}\left\lvert s-t\right\rvert^{2}.$ (19)
Observe that $\nabla V(s)=\int\nabla_{s}U(\theta,s)\pi(\mathrm{d}\theta\mid s)$. Substituting (18) and (19) into (17), we obtain
$\mathrm{KL}(\pi(\cdot\mid s),\pi(\cdot\mid t))$ $\leq\int\nabla_{s}U(\theta,s)^{T}(t-s)\pi(\mathrm{d}\theta\mid s)-\nabla V(s)^{T}(t-s)+\frac{L+L^{\prime}}{2}\left\lvert s-t\right\rvert^{2}$
$=\left(L+\frac{L^{2}}{2\mathtt{c}_{2}}\right)\left\lvert s-t\right\rvert^{2}.$
@@
$\leq\sqrt{\frac{L}{2}+\frac{L^{2}}{4\mathtt{c}_{2}}}\left\lvert s-t\right\rvert$.
@@
Consider a $L$-log-smooth target distribution $\pi$ that satisfies $\mathrm{PI}_{1}$ with constant $\mathtt{p}_{1}$. Suppose that for all $k\in[d]$ and $x\in\mathbb{R}^{d}$, the conditional distribution $\pi(\cdot\mid x_{-k})$ satisfies $\mathrm{PI}_{2}$ with uniform constant $\mathtt{c}_{2}$. Then, the conductance of the lazy systematic-scan Gibbs kernel is bounded by
$\Phi(\widetilde{P}_{\mathrm{SS}})\gtrsim\min\left\{1,\frac{\sqrt{\mathtt{c}_{2}}\mathtt{p}_{1}}{\sqrt{\mathtt{c}_{2}L+L^{2}}d}\right\}.$
@@
$\tau(\zeta,\widetilde{P}_{\mathrm{SS}})\lesssim\max\left\{1,\frac{(\mathtt{c}_{2}L+L^{2})d^{2}}{\mathtt{c}_{2}\mathtt{p}_{1}^{2}}\right\}\cdot\log\left(\frac{\sqrt{\Omega}}{\zeta}\right).$
However, the bounds in Corollary 4.7 appear to be loose when $\pi$ is a Gaussian distribution, as suggested by comparisons with Roberts and Sahu (1997).
@@
Consider a $L$-log-smooth target distribution $\pi$ that satisfies $\mathrm{PI}_{1}$ with constant $\mathtt{p}_{1}$. Suppose that for all $k\in[d]$ and $x\in\mathbb{R}^{d}$, the conditional distribution $\pi(\cdot\mid x_{-k})$ satisfies $\mathrm{PI}_{2}$ with uniform constant $\mathtt{c}_{2}$. Then, the conductance of the random-scan Gibbs kernel is bounded by
$\Phi(P_{\mathrm{RS}})\gtrsim\frac{1}{d\cdot\log d}\cdot\min\left\{1,\frac{\sqrt{\mathtt{c}_{2}}\mathtt{p}_{1}}{\sqrt{\mathtt{c}_{2}L+L^{2}}d\cdot\log d}\right\}.$
@@
$\lambda_{2}(P_{\mathrm{RS}})\gtrsim\frac{1}{d\cdot\log d}\cdot\min\left\{1,\frac{\mathtt{c}_{2}\mathtt{p}_{1}^{2}}{(\mathtt{c}_{2}L+L^{2})(d\cdot\log d)^{2}}\right\}.$
@@
$\tau(\zeta,P_{\mathrm{RS}})\lesssim d\cdot\log d\cdot\max\left\{1,\frac{(\mathtt{c}_{2}L+L^{2})(d\cdot\log d)^{2}}{\mathtt{c}_{2}\mathtt{p}_{1}^{2}}\right\}\cdot\log\left(\frac{\sqrt{\Omega}}{\zeta}\right).$
Unfortunately, the spectral gap bound in Corollary 4.9 is suboptimal when $\pi$ is strongly log-concave and $L$-log-smooth. In such cases, Ascolani et al. (2026a, Corollary 3.7) established that
@@
Furthermore, we believe that the lower bound on the spectral gap in Theorem 3.8 is not tight for the Gibbs sampler. This belief is supported by the case of the random walk on the $d$-dimensional hypercube, for which the conductance and spectral gap are of the same order, making the first bound in Theorem 3.8 loose in this setting. The spectral gap for this chain is computed, for example, in Diaconis (1988, Chapter 3C), while the conductance is computed in Diaconis and Stroock (1991, Example 3.2); see also Levin and Peres (2017, Examples 12.16 and 13.11) and Trevisan (2013, Chapter 5) for more recent surveys. The Gibbs sampler shares a similar local structure to the $d$-dimensional hypercube: each update resamples a single coordinate from its exact conditional, in analogy to the single-bit flips in the hypercube walk.
If the warm start parameter satisfies $\Omega=\mathcal{O}(1)$, Wadia (2024) proved an upper bound of $\mathcal{O}(L^{2}d^{7.5}(\log d)^{2}/\mathtt{p}_{1}^{4})$ on the mixing time of the lazy random-scan Gibbs sampler $\widetilde{P}_{\mathrm{RS}}$. However, a direct comparison of their results with Corollary 5.7 is not straightforward. This is due both to the laziness of the sampler and the fact that Wadia (2024) constrain the Gibbs sampler to a convex subset of $\mathbb{R}^{d}$, which leads them to analyze mixing via the more restrictive notion of $s$-conductance.
@@ 5.2 Non-log-concave examples
Spherical log-concavity without global log-concavity.
Consider the distribution
$\pi_{1}(\mathrm{d}x)\propto e^{-\left\lvert\left\lvert x\right\rvert-1\right\rvert}\mathrm{d}x.$
This distribution is log-Lipschitz continuous but not log-concave. Writing the potential function of $\pi_{1}$ as
[…]
with $u(t)=\left\lvert t-1\right\rvert$, which is convex for $t>0$, $\pi_{1}$ is nevertheless spherically log-concave and symmetric in the sense of Bobkov (2003); Bonnefont et al. (2016). It follows from their results that $\pi_{1}$ satisfies $\mathrm{PI}_{2}$.
@@
$\pi_{2}(\mathrm{d}x)\propto e^{-\left\lvert x\right\rvert+2\cos\left\lvert x\right\rvert}\mathrm{d}x.$
[…]
Gaussian mixture.
Let $m_{1},m_{2}\in\mathbb{R}^{d}$ with $m_{1}=-m_{2}=(1,1,\dots,1)$, and consider the two-component mixture of isotropic Gaussian distributions
$\pi_{3}(\mathrm{d}x)\propto\left(\frac{1}{2}\cdot e^{-\left\lvert x-m_{1}\right\rvert^{2}/2}+\frac{1}{2}\cdot e^{-\left\lvert x-m_{2}\right\rvert^{2}/2}\right)\mathrm{d}x.$
This distribution is log-smooth but not log-concave. Nevertheless, since each Gaussian distribution satisfies $\mathrm{LSI}_{2}$ (Gross, 1975), it follows from Schlichting (2019, Theorem 2) that $\pi_{3}$ also satisfies $\mathrm{LSI}_{2}$. Moreover, each conditional distribution of $\pi_{3}$ can be written as
$\pi_{3}(\mathrm{d}x_{k}\mid x_{-k})\propto\left(w(x_{-k})\cdot e^{-\left\lvert x_{k}-1\right\rvert^{2}/2}+(1-w(x_{-k}))\cdot e^{-\left\lvert x_{k}+1\right\rvert^{2}/2}\right)\mathrm{d}x_{k},$
where $w(x_{-k})\in[0,1]$. This is a two-component mixture of one-dimensional Gaussian distributions centered at $1$ and 1-1. As shown in Chafaï and Malrieu (2010, Corollary 4.5); Schlichting (2019, Corollary 1), the $L^{2}$-Poincaré constant of such mixtures remains uniformly bounded over all values of the mixing weight $w(x_{-k})$. Hence, the conditional distributions of $\pi_{3}$ satisfy a uniform $L^{2}$-Poincaré inequality.
@@ 6 Conclusion
We established explicit conductance bounds for systematic-scan and random-scan Gibbs samplers over a broad class of target distributions, namely those satisfying a Poincaré or log-Sobolev inequality and possessing sufficiently regular conditional distributions (Theorem 4.7 and Corollary 4.9). These bounds directly provide estimates on the mixing time in total variation distance and remain valid for either log-Lipschitz or log-smooth target distributions. In doing so, we extend prior results beyond the log-concave setting (Ascolani et al., 2026a; Wadia, 2024) and present several examples.
Our work also complements recent results showing that the spectral gaps and mixing times of systematic-scan and random-scan Gibbs samplers are closely related (Gaitonde and Mossel, 2026; Chlebicka et al., 2025). While our analysis focuses on coordinate-wise updates, our results naturally extend to blocked Gibbs samplers under corresponding block-smoothness conditions, allowing for joint updates of coordinate subsets. Moreover, our work also has implications for Metropolis-within-Gibbs sampling schemes; Ascolani et al. (2026b) examine how the conductance of generic coordinate-wise schemes compares to that of exact Gibbs sampling, while Qin et al. (2025) perform a similar analysis for the spectral gap. Their results, combined with our bounds, provide a framework for studying the convergence of broader coordinate-wise sampling methods.
Regarding dimension dependence, it is clear that our bounds on the spectral gap and mixing time are not tight, particularly for Gaussian targets (Amit, 1996; Roberts and Sahu, 1997) and, more generally, for log-smooth and strongly log-concave distributions (Ascolani et al., 2026a). We attribute this to the generality of our assumptions, which might not fully exploit the underlying structure of these distributions, and to the use of total variation distance in defining the mixing times. A natural direction for future research is therefore to refine these conductance bounds for both variants of the Gibbs sampler, particularly in terms of their dependence on the dimension and warmness parameter, as studied in Chen et al. (2020). One possibility is to study the mixing time of the Gibbs sampler under alternative probability metrics, such as appropriate Wasserstein distances, as in Wang and Wu (2014). Beyond the conductance approach, one could also explore recent techniques that may yield sharper spectral gap estimates for the Gibbs sampler, such as quasi-telescoping inequalities (also known as the spectral telescope), as developed in Qin and Wang (2024).